Grade 8General Physics

what would be the duration of the year if the distance between the earth and sun gets doubled?
1]1032
2]129
3]365
4]730days

mgayathri
9 Years agoGrade 8

1 Answer

Gaurav Sen
9 Years ago
The correct answer is 1032 days.By the formula , (T2/T1)^2 = (R2/R1)^3..Therefore , T2/365 = root (8) ..Therefore, T2 = 365 x 2.828 = 1032